Grade 11Algebra

If (a+1),3a,(4a+2) are in ap. Then 7th term of the series is.....?

Sejal Sinha
8 Years agoGrade 11

1 Answer

Parnav Bansal
8 Years ago
First term = a+1
Common difference =3a-a-1 = 2a-1
Seventh term =first term +(7-1) common difference
                         =a+1+6(2a-1)
                         = a+1+12a-6
                          =13a-5
7th term = 13a-5